# EvalTree: Profiling Language Model Weaknesses via Hierarchical Capability Trees

### v1

@@ 5.1 Low-Performance Identification Assessment
Low-Performance Identification Assessment compares how low LM performance is in weaknesses identified by different methods. We assess all weakness profiling methods on the MATH (Hendrycks et al., 2021b) and WildChat10K (a subset we curated from WildChat (Zhao et al., 2024a)) benchmarks and randomly split each benchmark into profiling/test sets (see Appendix C for more configuration details). We constrain the minimum weakness profile size to compare the average performance across identified weaknesses and constrain the minimum number of associated instances to compare overall performance on all associated instances. To visualize the comparisons, we plot two curves in Figure 3: one with the minimum profile size $M^{\prime}$ (ranging from 1 to 20) on the x-axis and $\min\{\sum_{w_{i}\in W_{\tau}}F(A(w_{i}))/|W_{\tau}|\mid\forall{\tau},|W_{\tau}|\geq M^{\prime}\}$ on the y-axis, and another with the minimum associated instance number $N^{\prime}$ (ranging from 1 to the test set size) on the x-axis and $\min\{F(S_{\tau})\mid\forall{\tau},|S_{\tau}|\geq N^{\prime}\}$ on the y-axis, where $S_{\tau}=\bigcup_{w_{i}\in W_{\tau}}A(w_{i})$. EvalTree consistently achieves the lowest curve, demonstrating its superior precision in capturing true weaknesses compared to other methods. See Appendix E.2 for qualitative analysis.
@@ 5.2 Ground-Truth Weakness Assessment
Ground-Truth Weakness Assessment compares how precisely and comprehensively different weakness profiling methods capture ground-truth weaknesses on synthetic LM evaluation results with appropriate description granularities. We manually curated 10 ground-truth weaknesses at various granularities for MATH and WildChat10K. For each benchmark, we generated three synthetic evaluation results by sampling with different hyperparameters that shape the probability distribution. For a given weakness profile, we compute the F1 score based on the overlap of associated instances to measure both precision and comprehensiveness relative to the ground-truth weakness profile $W^{*}$. We plot a curve with $M^{\prime}$ (ranging from 1 to 20) on the x-axis and the F1 score of $W_{\tau}$, where $|W_{\tau}|=M^{\prime}$ [^1] If multiple thresholds $\tau$ for EvalTree result in the same profile size, we select the lowest $\tau$. Note that the same profile size does not necessarily imply identical weakness profiles., on the y-axis. All curves are shown in Figure 4 and Appendix E.3.3. We observe that for most $M^{\prime}$, the F1 scores achieved by EvalTree surpass the highest F1 scores obtained by the other two methods. For additional details and analysis, see Appendix E.3.1 and E.3.2.
@@ 5.3 Extrinsic Assessment: Weakness-Guided Training Data Collection
Extrinsic Assessment compares how effectively weakness profiles from different methods guide targeted data collection to improve the evaluated LM; here, we conducted proof-of-concept experiments using a data-generation LM to generate (synthetic) data inputs (Kim et al., 2024) for data collection. The generic-capability-guided data collection strategy uses a description of the targeted benchmark’s overall capability as guidance. For each weakness profiling method, we have a corresponding data collection strategy that randomly samples an identified weakness (in the weakness profile generated by the method) as guidance for generating each data input. For context, we also included the result in which training data inputs were directly sampled from the profiling set; however, we emphasize that this strategy has an inherently unfair advantage due to its distributional match to the test set and is not a direct point of comparison in our proof-of-concept experiments, which focus on LM developers’ real-world practice of collecting new finetuning data.
We started with Llama 3.1 8B Instruct (Dubey et al., 2024) for MATH and DeepSeek-Coder-Base 6.7B (Guo et al., 2024) for DS-1000 (Lai et al., 2023), following configurations in Appendix C. When generating an input, we randomly sampled 5 inputs from the profiling set as in-context examples for the data-generation LM. We compared the performance of different LMs on the test set. For all data collection strategies, we collected the same amount of finetuning data inputs, with the output produced by separately feeding the input to the data-generation LM. Refer to Appendix E.4 for more details. The results in Figure 5 demonstrate that the LM trained on EvalTree-guided synthetic data significantly outperformed other LMs. Notably, the EvalTree-guided data collection strategy even slightly outperformed directly sampling data from the profiling set. Therefore, EvalTree can provide effective, targeted signals to guide data collection to improve LM performance.
@@ 5.4 LM Usage Cost Comparison
EvalTree also incurs significantly lower LM usage costs than other methods. When each method identifies 20 weaknesses on MATH, the LM usage costs of TextDiff and QualEval were approximately 20 and 8 times higher than EvalTree’s cost, respectively. This occurs because EvalTree’s LM usage cost remains constant regardless of $|W|$, whereas the costs of the others scale roughly linearly with $|W|$. See Appendix E.5 for further details.
5.5 Analysis on Threshold $\tau$ for Node Extraction
We analyze how the choice of $\tau$ influences the nodes extracted by the algorithm in Section 3.2. We examine the LM performance on all extracted nodes as $\tau$ varies, referred to as weakness/strength nodes, i.e., nodes extracted by the algorithm where the LM’s performance is significantly lower/higher than a given threshold $\tau$. To do this, we use the profiling set to build the capability tree and extract weakness/strength nodes with varying thresholds $\tau$. We locate the position of each instance in the test set on the capability tree by computing its capability embedding and then traversing from the root guided by the embedding. Specifically, at each non-leaf node, we predict the child cluster to which the instance belongs (by comparing its capability embedding with the K-Means clustering centers and picking the closest one), determining which child’s subtree to traverse into next; we call an instance that enters a weakness/strength node’s subtree a weakness/strength instance and study LM performance on all weakness/strength instances as $\tau$ varies.
@@ 6 Further Applications of EvalTree
Identifying Flaws in Chatbot Arena Evaluation. We give an application example by showing how EvalTree exposes flaws in the human-voter-based evaluation practice of Chatbot Arena (Chiang et al., 2024). We begin by using EvalTree to profile LM weaknesses on Chatbot Arena. To do this, we construct the capability tree for Chatbot Arena, where EvalTree ranks 64 LMs at each node by computing Elo scores based on human comparison pairs for instances linked to the node; it then identifies weaknesses of strong LMs like GPT-4 (OpenAI, 2023) by extracting nodes where their ranking is unexpectedly low. The weakness profile reveals surprising patterns, leading us to discover that the identified weakness may not stem from the LM itself but from flaws in the evaluation practice. For instance, at the node “Facilitating inclusive, ethical, and strategic communication and engagement across diverse and sensitive contexts,” LMs such as Zephyr-7B-$\beta$ (Tunstall et al., 2023) and Alpaca 13B (Taori et al., 2023) rank significantly higher than GPT-4 and Claude-2.1 (Anthropic, 2023). We observed that this node contains many user instructions with toxic requests, where human voters tended to prefer models that provide toxic responses over well-aligned models that refuse to answer; more quantitative analysis is provided in Appendix F. This shows that the evaluation practice of Chatbot Arena allows uncontrolled user preferences to diverge from the values of LM development, producing potentially unreliable evaluation results. Because even minor misaligned preferences can significantly change LM rankings (Zhao et al., 2024b; Huang et al., 2025; Min et al., 2025), the need for improved evaluation practices is pressing. In this example, EvalTree provides actionable insights for refining evaluation practices.
@@ 7 Future Work
Future work can enhance EvalTree in several ways. For example, capability tree construction can be improved by optimizing the tree structure and capability descriptions, making its dimensionality and granularity more controllable by humans, exploring model-dependent hierarchical structures, and extending it beyond language to other modalities, etc. Beyond direct enhancements, capability trees can also support a variety of potential applications. For example, they can help analyze LM evaluation results to tailor benchmarks to specific needs, to provide actionable insights into training data mixture, etc. By moving beyond aggregate metrics from existing evaluations, EvalTree enables a more comprehensive and interpretable analysis of LM performance across diverse capabilities, providing a useful foundation for future innovations in understanding and improving LM capabilities.
@@ Acknowledgments
We thank Zirui Cheng, Scott Geng, Joongwon Kim, Kyle Lo, Ian Magnusson, Sewon Min, Marco Tulio Ribeiro, Weijia Shi, Luca Soldaini, Ming Zhong, and Ruiqi Zhong for the insightful discussions. We thank Jacqueline He, Sandy Kaplan, Siting Li, Stella Li, Ben Newman, Rui Qiao, Rui Xin, and Lifan Yuan for proofreading the paper draft. We thank Hamish Ivison and Yuxuan Tong for sharing the model evaluation results. We thank members from the UW NLP and UW ML group for providing helpful feedback. We also thank All Hands AI’s product OpenHands (Wang et al., 2024b) and Xingyao Wang for their help with web interface implementation. This work is supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Information under the AI Visiting Professorship Programme (award number AIVP-2024-001), and by the AI2050 program at Schmidt Sciences.
@@ Appendix A Implementation Details of Automatic Capability Tree Construction
Recursive Clustering-Based Construction. As we mentioned in the main text above, clusterings are generated for each cluster number from 2 to a predefined maximum value, and the Silhouette score [^2] https://scikit-learn.org/1.5/modules/generated/sklearn.metrics.silhouette_score.html. All hyperparameters are set to their default values. (Rousseeuw, 1987), which measures clustering quality based on cohesion and separation, is computed for each clustering. In our experiments, the predefined maximum value is set to 10 by default. One detail is that, if no clustering achieves a positive score, all instances linked to the current node are treated as leaves and become direct children of it. For the K-Means implementation, we use sklearn.cluster.KMeans [^3] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html. All hyperparameters are set to their default values..
@@ Appendix B Implementation Details of Extracting Nodes with Low Performance
Algorithm 1 provides the pseudocode for extracting nodes with significantly low accuracy on the capability tree, the algorithm introduced in Section 3.2. In the pseudocode, we use Size to indicate the number of instances linked to a node.
@@
  Output: Set of extracted nodes $\mathcal{R}$
@@
  Node Extraction
@@ Appendix C Default Experimental Configurations
This section provides the experimental configurations used in Sections 5.15.3, and 5.5.
@@ E.2 Qualitative Analysis of Low-Performance Identification Assessment
Table 25 presents the identified weaknesses from TextDiff, QualEval, and EvalTree when the weakness profile size is 10, along with the LM performance on the associated instances (in the test set) of each identified weakness; they are based on applying the three methods to Llama 3.1 8B Instruct (Dubey et al., 2024) evaluation result on MATH (see Section 5.1). We observe that EvalTree-identified weakness descriptions are generally more specific than those identified by the other two methods, enabling a more precise diagnosis and thus capturing capabilities where the LM exhibits lowerer performance.
@@
Table 25: Weakness profiles generated by TextDiff, QualEval, and EvalTree, along with the LM performance on the associated instances (in the test set) of each identified weakness. Methods are run on Llama 3.1 8B Instruct (Dubey et al., 2024) evaluation result on MATH (Hendrycks et al., 2021b)
@@ E.3.1 Details of the Assessment Setup
This subsection provides additional details about the setup of Ground-Truth Weakness Assessment for weakness profiling methods in Section 5.2, based on the setup introduced in Section 2.2.
We used two benchmarks as testbeds, the MATH benchmark (Hendrycks et al., 2021b) and the WildChat10K benchmark (Zhao et al., 2024a). As described above, we manually curated a set of 10 ground-truth weaknesses (described in natural language) at diverse granularities as the ground-truth weakness profile, for MATH and WildChat10K, respectively. The ground-truth weakness profiles for MATH and WildChat10K are provided in Table 26 and Table 27, denoted as $W^{*}$. We aim to generate a synthetic evaluation result (on the profiling set) $g$ where the actual weaknesses are exactly this predefined weakness profile. First, we identify the associated instances for each ground-truth weakness. We then define two hyperparameters, the base probability $p\in(0,1]$ and the decrease rate $d\in(0,1)$, for controlling the sampling process. Taking correctness-based accuracy as an example, for the $i$-th benchmark instance, we compute the probability of it being solved correctly (i.e., $\mathbb{P}[g_{i}=1]$) as $p\times d^{m}$, where $m$ is the number of ground-truth weaknesses for which the instance is an associated instance. Finally, we independently sample correctness (1 or 0) for each $g_{i}$ using these computed probabilities, resulting in a synthetic evaluation result (on the profiling set). By design, the ground-truth weakness profile exactly represents the real weaknesses for this generated synthetic evaluation result, as we were mimicking the evaluation behavior of a hypothetical LM with exactly these weaknesses. As we described above, when using correctness-based accuracy as the metric for MATH, $p\times d^{m}$ represents the probability of an instance’s evaluation result being correct. Similarly, when using win-rate as the metric for WildChat10K, $p\times d^{m}$ denotes the probability of the (hypothetic) evaluated LM being preferred by the LM judge; specifically, we simulate the judge’s preference by sampling twice, once for the original order of responses and once after swapping their order (see Appendix C). For each benchmark, we generated three synthetic evaluation results using the hyperparameters $p=0.7$ and $d\in\{0.2,0.4,0.5\}$.

### v2

@@ 5.1 Low-Performance Identification Assessment
Low-Performance Identification Assessment compares how low the LM’s performance is on weaknesses identified by different methods. We assess all weakness profiling methods on the MATH (Hendrycks et al., 2021b) and WildChat10K (a subset we curated from WildChat (Zhao et al., 2024a)) benchmarks and randomly split each benchmark into profiling/test sets (see Appendix C for more configuration details). We constrain the minimum weakness profile size to compare the average performance across identified weaknesses and constrain the minimum number of associated instances to compare overall performance on all associated instances. To visualize the comparisons, we plot two curves in Figure 3: one with the minimum profile size $M^{\prime}$ (ranging from 1 to 20) on the x-axis and $\min\{\sum_{w_{i}\in W_{\tau}}F(A(w_{i}))/|W_{\tau}|\mid\forall{\tau},|W_{\tau}|\geq M^{\prime}\}$ on the y-axis, and another with the minimum associated instance number $N^{\prime}$ (ranging from 1 to the test set size) on the x-axis and $\min\{F(S_{\tau})\mid\forall{\tau},|S_{\tau}|\geq N^{\prime}\}$ on the y-axis, where $S_{\tau}=\bigcup_{w_{i}\in W_{\tau}}A(w_{i})$. EvalTree consistently achieves the lowest curve, demonstrating its superior precision in capturing true weaknesses compared to other methods. See Appendix E.2 for qualitative analysis.
@@ 5.2 Ground-Truth Weakness Assessment
Ground-Truth Weakness Assessment compares how precisely and comprehensively different weakness profiling methods capture ground-truth weaknesses (on synthetic LM evaluation results) with appropriate description granularities. We manually curated 10 ground-truth weaknesses at various granularities for MATH and WildChat10K. For each benchmark, we generated three synthetic evaluation results by sampling with different hyperparameters that shape the probability distribution. For a given weakness profile, we compute the F1 score based on the overlap of associated instances to measure both precision and comprehensiveness relative to the ground-truth weakness profile $W^{*}$. We plot a curve with $M^{\prime}$ (ranging from 1 to 20) on the x-axis and the F1 score of $W_{\tau}$, where $|W_{\tau}|=M^{\prime}$ [^2] If multiple thresholds $\tau$ for EvalTree result in the same profile size, we select the lowest $\tau$. Note that the same profile size does not necessarily imply identical weakness profiles., on the y-axis. All curves are shown in Figure 4 and Appendix E.3.3. We observe that for most $M^{\prime}$, the F1 scores achieved by EvalTree surpass the highest F1 scores obtained by the other two methods. For additional details and analysis, see Appendix E.3.1 and E.3.2.
@@ 5.3 Extrinsic Assessment: Weakness-Guided Training Data Collection
Extrinsic Assessment compares how effectively weakness profiles from different methods guide targeted training data collection to improve the evaluated LM; here, we conducted proof-of-concept experiments using a data-generation LM to generate (synthetic) data inputs (Kim et al., 2024) for data collection. The generic-capability-guided data collection strategy uses a description of the targeted benchmark’s overall capability as guidance. For each weakness profiling method, we have a corresponding data collection strategy that randomly samples an identified weakness (in the weakness profile generated by the method) as guidance for generating each data input. For context, we also included the result in which training data inputs were directly sampled from the profiling set; however, we emphasize that this strategy has an inherently unfair advantage due to its distributional match to the test set and is not a direct point of comparison in our proof-of-concept experiments, which focus on LM developers’ real-world practice of collecting new finetuning data.
We started with Llama 3.1 8B Instruct (Dubey et al., 2024) for MATH and DeepSeek-Coder-Base 6.7B (Guo et al., 2024) for DS-1000 (Lai et al., 2023), following configurations in Appendix C. When generating an input, we randomly sampled 5 inputs from the profiling set as in-context examples for the data-generation LM. We compared the performance of different LMs on the test set. For all data collection strategies, we collected the same amount of finetuning data inputs, with the output produced by separately feeding the input to the data-generation LM. Refer to Appendix E.4 for more details. The results in Figure 5 demonstrate that the LM trained on EvalTree-guided synthetic data significantly outperformed other LMs. Notably, the EvalTree-guided data collection strategy even slightly outperformed directly sampling data from the profiling set. Therefore, EvalTree provides effective and targeted signals for guiding data collection to improve LM performance.
@@ 5.4 LM Usage Cost Comparison
EvalTree also incurs significantly lower LM usage costs than other methods. When each method identifies 20 weaknesses on MATH, the LM usage costs of TextDiff and QualEval were approximately 20 and 8 times higher than EvalTree’s cost, respectively. This occurs because EvalTree’s LM usage cost remains constant regardless of the weakness profile size $|W|$, whereas the costs of the others scale linearly with $|W|$. See Appendix E.5 for details.
5.5 Analysis on Threshold $\tau$ for EvalTree’s Node Extraction
We analyze how the choice of $\tau$ influences the nodes extracted by the algorithm in Section 3.2. We examine the LM performance on all extracted nodes as $\tau$ varies, referred to as weakness/strength nodes, i.e., nodes extracted by the algorithm where the LM’s performance is significantly lower/higher than a given threshold $\tau$. To do this, we use the profiling set to build the capability tree and extract weakness/strength nodes with varying thresholds $\tau$. We locate the position of each instance in the test set on the capability tree by computing its capability embedding and then traversing from the root guided by the embedding. Specifically, at each non-leaf node, we predict the child cluster to which the instance belongs (by comparing its capability embedding with the K-Means clustering centers and then picking the closest one), determining which child’s subtree to traverse into next; we call an instance that enters a weakness/strength node’s subtree a weakness/strength instance and study LM performance on all weakness/strength instances from the test set as $\tau$ varies.
@@ 6 Further Applications of EvalTree
Identifying Flaws in Chatbot Arena Evaluation. We give an application example by showing how EvalTree exposes flaws in the human-voter-based evaluation practice of Chatbot Arena (Chiang et al., 2024). We begin by using EvalTree to profile LM weaknesses on Chatbot Arena. To do this, we construct the capability tree for Chatbot Arena, where EvalTree ranks 64 LMs at each node by computing Elo scores based on human comparison pairs for instances linked to the node; it then identifies weaknesses of strong LMs like GPT-4 (OpenAI, 2023) by extracting nodes where their ranking is unexpectedly low. The weakness profile reveals surprising patterns, leading us to discover that the identified weakness may not stem from the LM itself but from flaws in the evaluation practice. For instance, at the node “Facilitating inclusive, ethical, and strategic communication and engagement across diverse and sensitive contexts,” LMs such as Zephyr-7B-$\beta$ (Tunstall et al., 2023) and Alpaca 13B (Taori et al., 2023) rank significantly higher than GPT-4 and Claude 2.1 (Anthropic, 2023). We observed that this node contains many user instructions with toxic requests, where human voters tended to prefer models that provide toxic responses over well-aligned models that refuse to answer; more quantitative analysis is provided in Appendix F. This shows that the evaluation practice of Chatbot Arena allows uncontrolled user preferences to diverge from the values of LM development, producing potentially unreliable evaluation results. Because even minor misaligned preferences can significantly change LM rankings (Zhao et al., 2024b; Huang et al., 2025; Min et al., 2025), the need for improved evaluation practices is pressing. In this example, EvalTree provides actionable insights for refining evaluation practices.
@@ 7 Future Work
Future work can enhance EvalTree in several ways. For example, capability tree construction can be improved by optimizing the tree structure and capability descriptions, making its dimensionality and granularity more controllable by humans, exploring model-dependent hierarchical structures, and extending it beyond language to other modalities, etc. Additionally, it is useful to study how to quantitatively compare two capability trees directly. Beyond direct enhancements, capability trees can also support a variety of potential applications. For example, they can help analyze LM evaluation results to tailor benchmarks to specific needs, to provide actionable insights into training data mixture, etc. By moving beyond aggregate metrics from existing evaluations, EvalTree enables a more comprehensive and interpretable analysis of LM performance across diverse capabilities, providing a useful foundation for future innovations in understanding and improving LM capabilities.
@@ Acknowledgments
We thank Zirui Cheng, Scott Geng, Joongwon Kim, Kyle Lo, Ian Magnusson, Sewon Min, Marco Tulio Ribeiro, Weijia Shi, Luca Soldaini, Ming Zhong, and Ruiqi Zhong for the insightful discussions. We thank Jacqueline He, Sandy Kaplan, Siting Li, Stella Li, Jiacheng Liu, Ben Newman, Rui Qiao, Rui Xin, and Lifan Yuan for proofreading the paper draft. We thank Hamish Ivison and Yuxuan Tong for sharing the model evaluation results. We thank members from the UW NLP and UW ML group for providing helpful feedback. We also thank All Hands AI’s product OpenHands (Wang et al., 2024b) and Xingyao Wang for their help with web interface implementation. This work is supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Information under the AI Visiting Professorship Programme (award number AIVP-2024-001); by the AI2050 program at Schmidt Sciences; by a Google ML and Systems Junior Faculty Award; by NSF Grant Nos. IIS2142739 and IIS2044660; by the Defense Advanced Research Projects Agency’s (DARPA) SciFy program (Agreement No. HR00112520300); and by gift funding from Ai2.
@@ Appendix A Implementation Details of Automatic Capability Tree Construction
Recursive Clustering-Based Construction. As we mentioned in the main text above, clusterings are generated for each cluster number from 2 to a predefined maximum value, and the Silhouette score [^3] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.silhouette_score.html. All hyperparameters are set to their default values. (Rousseeuw, 1987), which measures clustering quality based on cohesion and separation, is computed for each clustering. In our experiments, the predefined maximum value is set to 10 by default. One detail is that, if no clustering achieves a positive score, all instances linked to the current node are treated as leaves and become direct children of it. For the K-Means implementation, we use sklearn.cluster.KMeans [^4] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html. All hyperparameters are set to their default values..
@@ Appendix B Implementation Details of Extracting Nodes with Low Performance
Algorithm 1 provides the pseudocode for extracting nodes with significantly low accuracy on the capability tree (the algorithm introduced in Section 3.2). In the pseudocode, we use Size to indicate the number of instances linked to a node.
@@
  Output: a set of extracted nodes $\mathcal{R}$
@@
  Second Pass: Node Extraction
@@ Appendix C Default Experimental Configurations
This section provides the experimental configurations used in Section 5.
@@ E.2 Qualitative Analysis of Low-Performance Identification Assessment
Table 25 presents the identified weaknesses from TextDiff, QualEval, and EvalTree when the weakness profile size is 10, along with the LM performance on the associated instances (in the test set) of each identified weakness; they are based on applying the three methods to Llama 3.1 8B Instruct (Dubey et al., 2024) evaluation result on MATH (see Section 5.1). We observe that EvalTree-identified weakness descriptions are generally more specific than those identified by the other two methods, enabling a more precise diagnosis and thus capturing capabilities where the LM exhibits lower performance.
@@
Table 25: Weakness profiles generated by TextDiff, QualEval, and EvalTree, along with the LM performance on the associated instances (in the test set) of each identified weakness. Methods are run on Llama 3.1 8B Instruct (Dubey et al., 2024) evaluation result on MATH (Hendrycks et al., 2021b).
@@ E.3.1 Details of the Assessment Setup
This subsection provides additional details about the setup of Ground-Truth Weakness Assessment in Section 5.2, based on the setup introduced in Section 2.2.
We used two benchmarks as testbeds, the MATH benchmark (Hendrycks et al., 2021b) and the WildChat10K benchmark (Zhao et al., 2024a). As described above, we manually curated a set of 10 ground-truth weaknesses (described in natural language) at diverse granularities as the ground-truth weakness profile, for MATH and WildChat10K, respectively. The ground-truth weakness profiles for MATH and WildChat10K are provided in Table 26 and Table 27, denoted as $W^{*}$. We aim to generate a synthetic evaluation result (on the profiling set) $g$ where the actual weaknesses are exactly this predefined ground-truth weakness profile $W^{*}$. First, we identify the associated instances for each ground-truth weakness. We then define two hyperparameters, the base probability $p\in(0,1]$ and the decrease rate $d\in(0,1)$, for controlling the sampling process. Taking correctness-based accuracy as an example, for the $i$-th benchmark instance, we compute the probability of it being solved correctly (i.e., $\mathbb{P}[g_{i}=1]$) as $p\times d^{m}$, where $m$ is the number of ground-truth weaknesses in $W^{*}$ for which the instance is an associated instance. Finally, we independently sample correctness (1 or 0) for each $g_{i}$ using these computed probabilities, resulting in a synthetic evaluation result (on the profiling set). By design, the ground-truth weakness profile $W^{*}$ exactly represents the real weaknesses for this generated synthetic evaluation result, as we were mimicking the evaluation behavior of a hypothetical LM with exactly these weaknesses. As we described above, when using correctness-based accuracy as the metric for MATH, $p\times d^{m}$ represents the probability of an instance’s evaluation result being correct. Similarly, when using win-rate as the metric for WildChat10K, $p\times d^{m}$ denotes the probability of the (hypothetic) evaluated LM being preferred by the LM judge; specifically, we simulate the judge’s preference by sampling twice, once for the original order of responses and once after swapping their order (see Appendix C). For each benchmark, we generated three synthetic evaluation results using the hyperparameters $p=0.7$ and $d\in\{0.2,0.4,0.5\}$.